OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) If one is added to the reciprocal of three times a number, the result is five times the reciprocal of the number. What is the number?

OpenStudy (anonymous):

\[\frac{1}{3x}+1=\frac{5}{x}\] would be my interpretation, but i could be wrong

OpenStudy (anonymous):

Can you help me with one more problem please? :) That was correct @satellite73

OpenStudy (anonymous):

Two cars leave town A at the same time and travel to town B, 480 miles away. The speed of the slower car is 20 mph less than the speed of the faster car. If the slower car arrives at town B two hours after the faster car, what is the speed of the slower car?

OpenStudy (anonymous):

i hate word problems

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

but we can do it anyways i am sure

OpenStudy (anonymous):

well any help would be amazing :)

OpenStudy (anonymous):

just gotta think of how to use distance is rate times time the rate of the slower car is what we are looking for so lets call it \(x\) that makes the rate of the faster car \(x+20\)

OpenStudy (anonymous):

we don't know the time, but we know \[x(T+2)=480\\ (x+20)T=480\] maybe we can solve this for \(x\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I hate these problems too...

OpenStudy (anonymous):

we are done really i think if we solve we will get \(x=60\) lets check the answer

OpenStudy (anonymous):

That's correct! Thank you soo much! :)

OpenStudy (anonymous):

the slower car going 60 mph will take \(480\div 60=8\) hours and the faster car going 80 mph will take (480/div 80=6\) hours